OpenStudy (anonymous):

HClO is a weak acid (Ka = 4.0 × 10–8) and so the salt NaClO acts as a weak base. What is the pH of a solution that is 0.059 M in NaClO at 25 °C?

OpenStudy (anonymous):

Na is a spectator ion in this situation. Start by finding the Kb of ClO–.

OpenStudy (anonymous):

help?

OpenStudy (zale101):

Before going ahead to do the ice diagram, you have to know that you are looking for the pH of a base. To do that, use the Henderson–Hasselbalch equation pOH=pKb+log[product]/[reactant] However, we have a ka not a kb. To get the Kb use this rule: \(\large Kw=Kb*Ka\) \(\large Kb=\frac{Kw}{Ka}\) Where Kw is 10^(-14)

OpenStudy (zale101):

That's solving for the Kb of ClO-

OpenStudy (zale101):

Now, to do the ice diagram, you have.